# Time Evidence Fusion Network: Multi-source View in Long-Term Time Series Forecasting

### v2

@@ Abstract.
In practical scenarios, time series forecasting necessitates timeliness, especially when dealing with large datasets. Consequently, the exploration of model architectures remains a perennially trending topic in research. To meet these performance demands, we propose a novel backbone from the perspective of information fusion. Introducing the Basic Probability Assignment (BPA) Module and the Time Evidence Fusion Network (TEFN), based on evidence theory, allows us to achieve superior performance. On the other hand, the perspective of multi-source information fusion effectively improves the accuracy of forecasting. Due to the fact that BPA is generated by fuzzy theory, TEFN also has considerable interpretability. In real data experiments, the TEFN partially achieved state-of-the-art, with low errors comparable to PatchTST, and operating efficiency surpass performance models such as Dlinear. Meanwhile, TEFN has high robustness and small error fluctuations in the random hyperparameter selection. TEFN is not a model that achieves the ultimate in single aspect, but a model that balances performance, accuracy, stability, and interpretability.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/ztxtech/Time-Evidence-Fusion-Network.
1. Introduction
The evolution of various phenomena is inherently tied to the progression of time, leading to an increasing prevalence of time series data in a multitude of application domains. From health diagnostics, where the progression of diseases is tracked over time, to financial analysis, where market trends are monitored for investment strategies, and weather forecasting, where atmospheric changes are predicted to safeguard communities, the importance of time series forecasting cannot be overstated. However, as the volume of data grows exponentially, time series forecasting faces significant performance challenges. The sheer scale of big data necessitates more efficient and faster algorithms to process and predict within acceptable time frames, which has become a pressing issue in the field.
@@ 1. Introduction
Figure 1. The channel dimension and time dimension of multivariate time series can be considered as different sources of information. Firstly, merge information sources of the same type, and finally integrate information sources from both channel and time dimensions to complete the final prediction.
In response to these challenges, researchers have been focusing on optimizing the structure of the backbone models used in time series forecasting. It has been observed that linear-based models, such as Dlinear (Zeng et al., 2023), often exhibit superior performance compared to transformer-based models like Crossformer (Zhang and Yan, 2023), particularly in terms of computational efficiency and speed. These findings highlight the potential of structural optimizations in enhancing the capabilities of time series forecasting models. Nevertheless, the quest for an ideal model that balances performance, prediction accuracy, stability, and interpretability continues. The aim is to develop a model that can provide robust support for high-stakes scenarios where timely and reliable predictions are critical, thereby bridging the gap between theoretical advancements and practical application needs.
Evidence theory, also known as Dempster-Shafer theory, provides a mathematical framework for representing and combining degrees of belief in the presence of uncertainty (Shafer, 1976; DEMPSTER, 1967). It is particularly useful in situations where information is incomplete, ambiguous, or comes from diverse sources. The theory is grounded in the concept of a belief function, which quantifies the degree to which a hypothesis is supported by the available evidence, without necessarily committing to a probability distribution over the possible outcomes. One of the key strengths of evidence theory is its ability to handle the fusion of multiple information sources effectively. This makes it a powerful tool in fields such as data fusion, sensor networks, and multi-criteria decision-making, where diverse pieces of information need to be integrated to form a coherent union. In the realm of symbolic logic, evidence theory often appears as a reasoning framework that offers high interpretability. Unlike traditional probability theory, which requires precise probability assignments, evidence theory allows for a more nuanced representation of uncertainty. It does this through the use of belief and plausibility measures, which provide a range of values for the likelihood of an event occurring, rather than a single probability. An innovative application of evidence theory in machine learning is the development of evidence neural networks by Denoeux (Denoeux, 2000). These networks are designed for classification tasks and integrate the principles of evidence theory into the architecture of neural networks. By doing so, they enhance the network’s ability to handle uncertainty and provide a more robust and interpretable decision-making process. The evidence neural networks output belief masses that can be directly interpreted as measures of evidence supporting each class, thus offering a clear understanding of the network’s reasoning process.
Therefore, we introduce the Time Evidence Fusion Network (TEFN), a novel architecture designed specifically for regression tasks within the realm of time series forecasting. The conceptual underpinning of TEFN is the innovative strategy of dissecting the time series data into dual perspectives—time and channels—thereby treating each channel and temporal instance as a distinct and autonomous source of information. This approach allows us to leverage the principles of evidence theory to fuse these diverse information streams within an integrated framework, aiming to enhance the accuracy and reliability of predictions as depicted in Figure 1. In essence, TEFN is built upon the premise of simplicity and naivety in its design philosophy. It refrains from incorporating complex technical modifications that often render models opaque and difficult to interpret. Instead, TEFN maintains a straightforward structure that prioritizes clarity and ease of understanding, making it a transparent and accessible backbone for time series analysis. By adopting this minimalist approach, TEFN emerges as a refreshing alternative in the field, offering a new foundation upon which more sophisticated forecasting models can potentially be built, while retaining the interpretability and trustworthiness that are hallmarks of evidence theory.
In the architecture of TEFN, we introduce a pivotal component termed the Basic Probability Assignment (BPA) Module, which stands as the core of our evidence-theoretic approach to time series forecasting. The BPA Module operates on the fundamental principles of evidence theory, meticulously mapping various sources of information into mass distributions that are intrinsically linked to the desired target outcome. This mapping is achieved through the application of fuzzy membership functions, which assign membership grades to different fuzzy sets, thereby constructing a comprehensive target distribution as proposed by Deng (Deng, 2015). In contrast to the convolutional operations prevalent in image processing, which selectively condense features pertinent to the target through filters (Bengio et al., 1993), the BPA Module embodies an expansionary process, as illustrated in Figure 2. Rather than narrowing down to specific features, BPA embraces an expansive perspective, contemplating all potentialities within the event space, which is a collection of subsets derived from the sample space. This expansion across various events is particularly advantageous for the simplistic structures inherent in time series data, enabling a thorough exploration of the myriad hidden informational layers within. For the diverse distributions yielded by the multitude of information sources, we treat each source as an individual sample, and by amalgamating these samples, we utilize the expected forecasting values as the ultimate prediction, thereby delivering a robust and informed forecast that encapsulates the collective wisdom of all contributing sources.
@@
Figure 2. Comparison between convolution and basic probability assignment: convolution condenses information while basic probability assignment diverges to consider different possibilities.
In this paper, we delve into the practical nuances of the Time Evidence Fusion Network (TEFN), examining its precision, memory consumption, training velocity, sensitivity to parameters, and interpretability. To maintain equitable comparisons, all experimental data and configuration files are sourced from open-source initiatives. TEFN has partially achieved state-of-the-art (SOTA) performance on various real-world datasets, with a notable reduction in training duration and model complexity when contrasted with Transformer-based models such as Crossformer (Zhang and Yan, 2023) and Patch-based models like PatchTST (Nie et al., 2023). In terms of random hyperparameter selection, TEFN demonstrates minimal variability, affirming its stability. The contributions of this manuscript are summarized as follows:
  • The proposed method is well-suited for large datasets, showcasing exceptional performance.
@@
2. Related Work
2.1. Dempster-Shafer Theory
Dempster-Shafer theory, also known as Evidence Theory, operates on a different framework compared to traditional probability theory (Shafer, 1976; DEMPSTER, 1967). It relies on the concept of mass functions, which provide a more flexible representation of uncertainty, accommodating weaker constraints than traditional probabilities. When combining two mass functions, denoted as $m_{1}$ and $m_{2}$, the resulting fused mass function $m$ is determined using the Dempster-Shafer Rule (DSR), as described in Equation 1. The symbols $A$, $B$, and $C$ represent focal elements, which are typically combinations of all independent samples (or target labels, such as the classification targets).
(1) $m(A)=\sum_{B\cap C=A}m_{1}(B)*m_{2}(C)$
2.2. Basic Probability Assignment
BPA is a core concept in evidence theory, used to map real data $x$ into mass distributions. Mass distributions describe the degree of support for various hypotheses in the event space, providing a basis for reasoning and decision-making. The generation of BPA $m$ is often based on fuzzy logic, where fuzzy membership functions $\mu$ are used to map data points to different fuzzy sets, thereby constructing the target distribution in Equation 2 where $w$ and $b$ is weight and bias of the membership function. BPA can also be generated by assigning support degrees through Gaussian distributions $N$ (probability density function $f$) where $\mu$ and $\sigma$ are expectation and variance in Equation 3.
(2) $\begin{split}m=\mu(x)=wx+b\end{split}$
(3) $\begin{split}&x\sim N(\mu,\sigma)\\ &m=f(x)=\frac{1}{\sqrt{2\pi\sigma^{2}}}e^{-\frac{(x-\mu)^{2}}{2\sigma^{2}}}\\ \end{split}$
@@ 2.2. Basic Probability Assignment
2.3. Evidence Decision Making
The process of Evidence Decision Making can be delineated into several sequential steps: basic probability assignment (BPA), evidence fusion, and decision making (DM) (Deng, 2015; Zhan et al., 2024). BPA serves as a pivotal initial phase, constituting a mapping from raw data to mass distribution, akin to the path from raw data to probability distribution. Subsequently, evidence fusion integrates distinct mass distributions from various data channels, culminating in a rationalized mass distribution. The most classic fusion method is the DSR (Shafer, 1976; DEMPSTER, 1967) mentioned in Section 2.1. Finally, decision making transpires as the progression from the attained distribution towards the designated target. The most classic decision-making method is to transform the mass distribution into the probability distribution of the target, called pignostic probability transformation (PPT) (Dubois et al., 1993) in Equation 4, where $x$ is a target label and $e$ are combinations of all possible target labels.
(4) $p(x)=\sum_{x\in e}\frac{m(e)}{(1-m(\emptyset)|e|)}$
Figure 3. The overall structure of TEFN
2.4. Linear Time Series Model
There are many linear time series forecasting models (LTSFM). LTSFM offer several advantages for modeling and forecasting time series data. Firstly, their simplicity and transparency make them easy to implement, interpret, and analyze. The parameters of linear models directly represent the relationships between input features and the output, providing insights into the underlying dynamics of the time series. Secondly, LTSFM exhibit high efficiency, requiring fewer computational resources and hyperparameters compared to complex models like transformer-based models (Vaswani, 2017; Zhang and Yan, 2023). This leads to faster training and inference times, making them suitable for real-time applications and large datasets.
A classic LTSFM model is Dlinear (Zeng et al., 2023). DLinear serves as a simple yet effective baseline for long-term time series forecasting. It is a model that employs a seasonal-trend decomposition technique to separate the trend and seasonal components of the time series data. Each component is then independently modeled using a single-layer linear regression. The final prediction is obtained by summing the outputs of the two linear regressions.
@@ 2.4. Linear Time Series Model
3. Time Evidence Fusion Network
The structure of TEFN is inherently straightforward, achieving remarkable performance with a design featuring only two BPA modules, akin to employing just two convolutional kernels. We posit that the novel architecture of TEFN embodies the essence of a pure backbone model, stripped of domain-specific intricacies. This minimalist approach allows TEFN to focus solely on capturing the fundamental temporal patterns within the data, making it a versatile and adaptable foundation for various time series analysis tasks. The overall structure of TEFN is depicted in Figure 3.
Figure 4. The operation process of a BPA module for a time series $T$ of length $L=5$. The sample space size in the BPA module is set as $|S|=2$, and the event space size is set as $|2^{S}|=2^{|S|}=4$. The operation of the BPA module is similar to that of the convolution module, but the BPA module diffuses different masses instead of summing them in the convolution.
3.1. Time Normalization and De-normalization
To expedite convergence towards local optima and enhance training efficiency, TEFN incorporates a normalization technique inspired by the Stationary model (Liu et al., 2022). This process computes the mean $\mu$ as shown in Equation 5 and variance $\sigma^{2}$ as shown in Equation 6 for each segment of the time series $x$. The subsequent normalization and de-normalization operations, presented in Equation 7 and Equation 8, respectively, act as symmetrical transformations that standardize the time series data. The neural network operation $Net(\cdot)$ is then applied to the normalized time series $x_{Norm}$, ensuring that the model learns from data with consistent scale and distribution. This normalization strategy effectively mitigates the influence of outliers and initial data scale, fostering faster convergence and improved stability during training.
(5) $\mu=\frac{1}{|x|}\sum_{x_{i}\in x}x_{i}$
(6) $\sigma^{2}=\frac{1}{|x|}\sum_{x_{i}\in x}(x_{i}-\mu)^{2}=\frac{1}{|x|}\sum_{x_{i}\in x}x_{i}^{2}-\mu^{2}$
(7) $x_{Norm}=\frac{x-\mu}{\sigma}$
(8) $x_{De-Norm}=\sigma*\hat{y}_{Norm}+\mu=\sigma*Net(x_{Norm})+\mu$
3.2. Time Dimension Projection
In TEFN, we employ a linear layer $project(\cdot)$ to project dimensionality to an expansion dimensionality, transforming the input time series $x$ with a length of $L_{in}$ into a sequence $x^{\prime}$ with a length of $L_{in}+L_{pred}$ in Equation 9. Here, $L_{pred}$ represents the desired length of the prediction horizon. This operation is analogous to the time projection mechanism utilized in other neural network architectures, such as TimesNet (Wu et al., 2023).
(9) $x^{\prime}=project(x)=x\times W_{p}+b$
@@ 3.2. Time Dimension Projection
3.3. Basic Probability Assignment Module
In TEFN, we leverage the principles of Evidence Theory to model the uncertainty and ambiguity inherent in time series data. According to Evidence Theory, each dimension of the time series can be represented as a mass function within a common event space. This event space is typically defined as the power set $2^{S}$ of the sample space $S$, where $S$ represents the possible values or labels for the data. For classification problems, $S$ corresponds to all the possible labels, while for regression problems, $S$ is treated as a set of anonymous samples, ensuring mutual exclusivity among them.
This representation allows us to capture the fuzzy nature of the data, where each dimension generates a mass distribution within the event space. To achieve this, we utilize parameterized fuzzy membership functions, as defined in Equation 10. Here, $D$ represents the dimensionality of the time series, encompassing both the time dimension $T$ and the channel dimension $C$. The indices $i$ and $j$ correspond to the specific dimensions within $D$, while $k$ represents the index of the event space dimension, denoted as $F$. The fuzzy membership function $\mu_{k}$ is defined in Equation 10 where $w_{D,j,k}$ and $b_{D,j,k}$ are the slope and intercept parameters of the membership function, respectively. These parameters are learned during training and allow the model to capture the varying degrees of membership for each data point within the event space.
(10) $m_{D,i,j,k}=\mu(x_{Norm,i,j})=w_{D,j,k}*x_{Norm,i,j}+b_{D,j,k}$

### v3

@@ Abstract
In practical scenarios, time series forecasting necessitates not only accuracy but also efficiency. Consequently, the exploration of model architectures remains a perennially trending topic in research. To address these challenges, we propose a novel backbone architecture named Time Evidence Fusion Network (TEFN) from the perspective of information fusion. Specifically, we introduce the Basic Probability Assignment (BPA) Module based on evidence theory to capture the uncertainty of multivariate time series data from both channel and time dimensions. Additionally, we develop a novel multi-source information fusion method to effectively integrate the two distinct dimensions from BPA output, leading to improved forecasting accuracy. Lastly, we conduct extensive experiments to demonstrate that TEFN achieves performance comparable to state-of-the-art methods while maintaining significantly lower complexity and reduced training time. Also, our experiments show that TEFN exhibits high robustness, with minimal error fluctuations during hyperparameter selection. Furthermore, due to the fact that BPA is derived from fuzzy theory, TEFN offers a high degree of interpretability. Therefore, the proposed TEFN balances accuracy, efficiency, stability, and interpretability, making it a desirable solution for time series forecasting.
Index Terms:
Time Series, Forecasting, Evidence Theory, Information Fusion, Neural Network
publicationid: pubid: $0$–$00$ 2021 IEEE
1 Introduction
The evolution of various phenomena is inherently tied to the progression of time, leading to an increasing prevalence of time series data in a multitude of application domains [1], such as medical diagnosis [2, 3, 4], education [5], weather [6, 7], power system [8], and so on [9, 10, 11]. The importance of time series forecasting cannot be overstated. There are also many methods for time series forecasting. The initial methods came from statistics, such as Autoregressive (AR) model, Moving Average (MA) model, Autoregressive Moving Average (ARMA) model, and Autoregressive integrated moving average (ARIMA) model [12]. Deep learning is a powerful method with wide applications in various disciplines [13, 14]. In recent years, time series prediction models based on neural networks have become a research hotspot. These neural networks have evolved from the simplest deep neural networks (DNN) [15]. Until now, there are already many backbones that can be used for time series, such as Recurrent Neural Network (RNN) [16], Long Short-Term Memory (LSTM) [17], Transformer [18, 19, 20, 21], and so on. Large Language Models (LLM) for time series also perform well in cross dataset forecasting tasks [22, 23, 24].
There is uncertainty in time series data, and modeling uncertainty on time series is effective. Evidence theory is an uncertain reasoning framework based on mass function [25, 26, 27, 28]. Previous studies have shown that evidence theory can effectively utilize the uncertainty of time-series data and achieve positive application results [29, 30]. On the other hand, evidence theory is also a method of information fusion that can effectively aggregate different information sources and make effective decisions [31, 32]. Meanwhile, evidence theory can also be applied in neural networks to accomplish tasks such as classification [33]. Therefore, in order to reduce the error of time series forecasting, we propose the Time Evidence Fusion Network (TEFN) from the perspectives of evidence theory and information fusion. TEFN is a lightweight neural network that considers the time and channel dimensions of multivariate time series as different sources of information in Figure 1.
@@ 1 Introduction
Figure 1: The channel dimension and time dimension of multivariate time series can be considered as different sources of information. Firstly, merge information sources of the same type, and finally integrate information sources from both channel and time dimensions to complete the final prediction.
In TEFN, we propose a novel Basic Probability Assignment (BPA) Module. Based on the evidence theory of decision-making process, BPA maps different sources of information to distributions related to the target outcome. The target distribution is achieved by the fuzzy membership of different fuzzy sets [34]. In image processing, convolution condenses the required features of the target [35]. Unlike convolution, BPA is an expansion process in Figure 2. On the contrary, BPA considers all possibilities in the event space composed of subsets of the sample space, expanding on different events, which is more effective for simple structures like time series and can fully explore the different hidden information inside. For different distributions generated by multiple information sources, we consider sources as samples and use the expected forecasting values as the final result.
@@
Figure 2: Comparison between convolution and basic probability assignment: convolution condenses information while basic probability assignment diverges to consider different possibilities.
In this article, we focus on the practical application details of TEFN, including its accuracy, memory usage, training speed, parameter sensitivity, and interpretability. At the same time, in order to ensure fairness, the experimental data and configuration files for comparison are all from open-source projects. TEFN achieves state-of-the-art (SOTA) on multiple real datasets, while training time and model parameters are much smaller than models based on Transformer [20, 36]. For random hyperparameter selection, TEFN exhibits weak fluctuations and is a stable model. The contributions of this article are as follows:
@@
2 Related Work
2.1 Dempster-Shafer Theory
Dempster-Shafer theory, also known as Evidence Theory, operates on a different framework compared to traditional probability theory [25, 26]. It relies on the concept of mass functions, which provide a more flexible representation of uncertainty, accommodating weaker constraints than traditional probabilities. All independent and mutually exclusive sample sets that exist in the system are called frames of discrete (FOD), represented as $\Omega$.
When there are multiple sources of information, information fusion is usually required to make decisions. Fusing two mass functions, denoted as $m_{1}$ and $m_{2}$, the fused mass function $m$ is determined using the Dempster-Shafer Rule (DSR), as described in Equation (1). The symbol variables $A$, $B$, and $C$ are focal elements, which are typically combinations of all independent samples in the FOD, which are often composed of target labels, such as the classification targets.
$m(A)=\sum_{B\cap C=A}m_{1}(B)*m_{2}(C)$ (1)
2.2 Basic Probability Assignment
BPA is a core concept in evidence theory, used to map real data $x$ into mass distributions. Mass distributions describe the degree of support for various hypotheses in the event space, providing a basis for reasoning and decision-making. The generation of BPA $m$ is often based on fuzzy logic, where fuzzy membership functions $\mu$ are used to map data points to different fuzzy sets, thereby constructing the target distribution in Equation (2) where $w$ and $b$ is weight and bias of the membership function. BPA can also be generated by assigning support degrees through Gaussian distributions $N$ (probability density function $f$) where $\mu$ and $\sigma$ are expectation and variance in Equation (3).
$\begin{split}m=\mu(x)=wx+b\end{split}$ (2)
$\begin{split}&x\sim N(\mu,\sigma)\\ &m=f(x)=\frac{1}{\sqrt{2\pi\sigma^{2}}}e^{-\frac{(x-\mu)^{2}}{2\sigma^{2}}}\\ \end{split}$ (3)
@@ 2.2 Basic Probability Assignment
2.3 Evidence Decision Making
The process of Evidence Decision Making can be delineated into several sequential steps: basic probability assignment (BPA), evidence fusion, and decision making (DM) [34, 37]. BPA serves as a pivotal initial phase, constituting a mapping from raw data to mass distribution, akin to the path from raw data to probability distribution. Subsequently, evidence fusion integrates distinct mass distributions from various data channels, culminating in a rationalized mass distribution. The most classic fusion method is the DSR [25, 26] mentioned in Section 2.1. Finally, decision making transpires as the progression from the attained distribution towards the designated target. The most classic decision-making method is to transform the mass distribution into the probability distribution of the target, called pignistic probability transformation (PPT) [38] in Equation (4), where $x$ is a target label and $e$ are combinations of all possible target labels.
$p(x)=\sum_{x\in e}\frac{m(e)}{(1-m(\emptyset)|e|)}$ (4)
2.4 Linear Time Series Model
There are many linear time series forecasting models (LTSFM) [12, 39, 40, 41]. LTSFM offer several advantages for modeling and forecasting time series data:
  • The simplicity and transparency make them easy to implement, interpret, and analyze.
  • This leads to faster training and inference times, making them suitable for real-time applications and large datasets.
  • No additional encoding layers are required.
A classic LTSFM model is Dlinear [39]. DLinear serves as a simple yet effective baseline for long-term time series forecasting. It is a model that employs a seasonal-trend decomposition technique to separate the trend and seasonal components of the time series data. Each component is then independently modeled using a single-layer linear regression. The final prediction is obtained by summing the outputs of the two linear regressions.
@@ 2.4 Linear Time Series Model
3 Time Evidence Fusion Network
The structure of TEFN is inherently straightforward, achieving remarkable performance with a design featuring only two BPA modules, akin to employing just two convolutional kernels. We posit that the novel architecture of TEFN embodies the essence of a pure backbone model, stripped of domain-specific intricacies. This minimalist approach allows TEFN to focus solely on capturing the fundamental temporal patterns within the data, making it a versatile and adaptable foundation for various time series analysis tasks. The overall structure of TEFN is depicted in Figure 3. Next, we will elaborate on each module one by one.
Figure 3: The overall structure of TEFN: Establish corresponding mass functions for the time dimension and channel dimension as the target dimensions of BPA.
Figure 4: The operation of a BPA module: weight matrix and bias matrix are both learnable parameters. BPA elevates the dimensionality of data by adding a dimension to distinguish different events.
3.1 Time Normalization and De-normalization
To expedite convergence towards local optima and enhance training efficiency, TEFN incorporates a normalization technique inspired by the Stationary model [42]. This process computes the mean $\mu$ as shown in Equation (5) and variance $\sigma^{2}$ as shown in Equation (6) for each segment of the time series $x$. The subsequent normalization and de-normalization operations, presented in Equation (7) and Equation (8), where $\hat{y}$ is forecasting result, respectively, act as symmetrical transformations that standardize the time series data. The neural network operation $Net(\cdot)$ is then applied to the normalized time series $x_{Norm}$, ensuring that the model learns from data with consistent scale and distribution. This normalization strategy effectively mitigates the influence of outliers and initial data scale, fostering faster convergence and improved stability during training.
$\mu=\frac{1}{|x|}\sum_{x_{i}\in x}x_{i}$ (5)
$\sigma^{2}=\frac{1}{|x|}\sum_{x_{i}\in x}(x_{i}-\mu)^{2}=\frac{1}{|x|}\sum_{x_{i}\in x}x_{i}^{2}-\mu^{2}$ (6)
$x_{Norm}=\frac{x-\mu}{\sigma}$ (7)
$\hat{y}=\sigma*\hat{y}_{Norm}+\mu=\sigma*Net(x_{Norm})+\mu$ (8)
3.2 Time Dimension Projection
In TEFN, we employ a linear layer $project(\cdot)$ to project dimensionality to an expansion dimensionality, transforming the input time series $x$ with a length of $L_{in}$ into a sequence $x^{\prime}$ with a length of $L_{in}+L_{pred}$ in Equation (9). Here, $L_{pred}$ represents the desired length of the prediction horizon. This operation is analogous to the time projection mechanism utilized in other neural network architectures, such as TimesNet [43].
$x^{\prime}=project(x)=x\times W_{p}+b$ (9)
@@ 3.2 Time Dimension Projection
3.3 Basic Probability Assignment Module
In TEFN, we leverage the principles of Evidence Theory to model the uncertainty and ambiguity inherent in time series data. According to Evidence Theory, each dimension of the time series can be represented as a mass function within a common event space. This event space is typically defined as the power set $2^{S}$ of the sample space $S$, where $S$ represents the possible values or labels for the data. Since time series forecasting is a regression problem, the size of the sample space needs to be used as a hyperparameter.
This representation allows us to capture the fuzzy features of the input time series, where each dimension generates a mass distribution within the event space. To achieve this, we utilize parameterized fuzzy membership functions, as defined in Equation (10). Here, $D$ represents the dimensionality of the time series, encompassing both the time dimension $T$ and the channel dimension $C$. The indices $i$ and $j$ correspond to the specific dimensions within $D$, while $k$ represents the index of the event space dimension, denoted as $F$. The fuzzy membership function $\mu_{k}$ is defined in Equation (10) where $w_{D,j,k}$ and $b_{D,j,k}$ are the slope and intercept parameters of the membership function. These parameters are learned during training and allow the model to capture the varying degrees of membership for each data point within the event space. Figure 4 shows the visible process of BPA module.
$m_{D,i,j,k}=\mu(x_{Norm,i,j})=w_{D,j,k}*x_{Norm,i,j}+b_{D,j,k}$ (10)